QUESTION 40

  1. There are two states of the world. In state 1, you get $100. In state 2, you get $50. Assuming rationality, and the expected return is $87.50, what is the probability state 1 occurs?

    0.25

    0.5

    0.75

    1
Answer #1

Answer

Option 3

the formula is: (note: the two events are mutually exhaustive and exclusive so the sum of probability of both is equal to 1)

E(2) = X* P(2) P(state 1) =1- P(state 2) P(state 1) = a P(state b) = 1-a 87.5 = 100 * a + 50 * (1 – a) = 100a + 50 – 50a = 50

the probability of state a=0.75